# Real $K$-Theory for $C^{*}$-Algebras: just the facts October 22, 2025

### v2

@@ Abstract.
This work is intended to present the basic properties of $KO$-theory for real $C^{*}$-algebras and to explain its relationship with complex $K$-theory and with $KR$- theory. Whenever possible we will rely upon proofs in printed literature, particularly the work of Karoubi [52], Wood [89], Schröder [76], and more recent work of Boersema [15] and J. M. Rosenberg [70]. In addition, we shall explain how $KO$-theory is related to the Ten-Fold Way in physics and point out how some deeper features of $KO$-theory for operator algebras may provide powerful new tools there. Commutative real $C^{*}$-algebras not of the form $\real(X)$ will play a special role.
Unfortunately, there is no single reference for $KO$-theory for operator algebras that begins to compare with Blackadar’s wonderful exposition of complex $K$-theory [14]. This work is intended to provide a platform upon which mathematicians and mathematical physicists can rely in order to use these new tools in their research. As we are writing for a diverse audience of functional analysts, topologists, and physicists, we often present material well-known to one group of people and unfamiliar to another.
@@ 1. Introduction
First, a little history to put $KO$-theory into historical context. The group $K_{0}(R)$ was introduced by Alexander Grothendieck in 1957 as part of his revolution in algebraic geometry, as a way to formulate what is now known as the Grothendieck-Riemann-Roch theorem. It was based upon coherent sheaves on an algebraic variety. It was extended using modules over arbitrary rings and is difficult to compute in general. [^1] For everything you would want to know about the early history of algebraic $K$-theory, see Weibel [87]. The group $K_{1}(R)$ was defined in 1962, $K_{2}(R)$ in 1967, and $K_{n}(R)$ in 1970.
@@
As their paper was about to appear, Atiyah’s student R. Wood produced (1966)[89] a proof of the Periodicity Theorem along the lines suggested by ABS, where Clifford algebras play an essential role and where Bott Periodicity is a consequence of eight homotopy equivalences. He took a fundamental step forward, in that he stated the theorem in terms of the general linear groups of real and complex Banach algebras, and not just compact spaces, though his proof was rather sketchy. (See also Milnor [62]). This was followed by several detailed papers by Max Karoubi (cf. [45]). Karoubi generalized the Periodicity Theorem to Banach categories and Banach rings, proving periodicity in detail. Karoubi and Orlando Villamayor ([50], [51] ) explored its relationship to algebraic $K$-theory. A few years later Karoubi introduced Hermitian $K$-theory ([47], [49]), where he proved what he calls “A Fundamental Theorem” which is another road to Bott Periodicity (see the introduction to [47]). This theorem has important consequences nowadays in algebraic geometry and in quadratic forms over any field, not just the field of real numbers. Karoubi anticipated many of the results that we use in this work , In particular, the idea of desuspension was used to produce the negative algebraic $K$-theory groups developed by Bass.
@@
This work is designed to provide a basic source for those in mathematics and physics who wish to use $KO$-theory in their work. There is no such source at present: we mean to remedy that problem.
@@
Section 6 In Section 6 we introduce the homology theory $KO_{*}$ for real $C^{*}$-algebras, the main focus of this work . There are several distinct approaches to $KO$. We use what we believe to be the simplest approach, not relating much to algebraic $K$-theory, Hermitian $K$-theory, or other homotopy constructions. We show that $KO_{*}$ is a homology theory and introduce some of its relationships to $K_{*}$. Still no mention of Bott periodicity.
@@
Section 9 Section 9 introduces basic constructions in topological spaces such as the product, the join, the smash product and the suspension, and establishes the fundamentally important identification $F(SX,Y)\cong F(X,\Omega Y)$. We will assume throughout the work that all topological spaces are compactly generated. We introduce $CW$-complexes briefly.
@@
Section 15 Section 15 is devoted to establishing the relationship between $KO^{*}$-theory for spaces and $KO_{*}$-theory for real $C^{*}$-algebras and similarly for $K$-theory and complex $C^{*}$-algebras. We present a novel proof of Swan’s classical result relating the $KO_{*}$-theory of a real $C^{*}$-algebra $C^{\mathbb{R}}(X)$ and the topological $KO^{*}$-theory of the compact space $X$.
@@
Section 24 This section presents a novel approach to the groups $KO_{*}(A)$, showing that each of the eight groups may be described concretely in terms of unitary elements of A and associated algebras such as $M_{k}(A)$.
Section 25 This section contains our treatment of $KO$-theory with coefficients. It is well-known that $H^{*}(X;G)$, the cohomology of a space with coefficients in a groups $G$ is sometimes much more useful and certainly more tractable then when you take $G={\mathbb{Z}}$. We suspect that the same will be true for $KO$ and hence we present how to define it.
What’s not in the work ? We have not discussed the Künneth theorem in full generality, nor united $K$-theory, which combines $K$, KOKO and $KSc$ theories. We have not discussed some of the applications of $KO$-theory in differential geometry, as this requires quite a bit of background to appreciate. Nor have we discussed $KO$-homology theory for compact spaces. We have not discussed the Hopkins-Hovey isomorphism [43]
@@
This work arose in connection with the second author’s collaboration with Erik Akkermanns. We are most grateful to him and to his students, particularly to Nadav Orion, who have provided invaluable feedback and true insight in the development of this work. We are also very grateful to Jean Bellissard, Joachim Cuntz, Yigal Kamel, John Klein, Chris Phillips, Emil Prodan, Jonathan Rosenberg and Chuck Weibel, for providing us with various deep insights as this work developed. We especially want to acknowledge with great pleasure the extensive philosophical and detailed help that we have received from Max Karoubi, who, more than anyone, is responsible for the very existence of $KO$-theory for spaces and algebras as we know it.
@@ 2. Real $C^{*}$-algebras
Two points on notation. First, we have been confused over the years by various papers using $C(X)$, $K_{0}(A)$, and the like, without specifying whether they had real or complex versions in mind.777and we each plead guilty to this sin. Frequently authors deal with this by stating at the beginning that everything is complex or that everything is real. In this work both possibilities will arise. We will be more precise.
A second point. Usually the real numbers occur in this work as a real vector space or with their real $C^{*}$-algebra structure, and when that happens we shall denote them as $\mathbb{R}$}. Sometimes, though, we want to think of the real numbers simply as a locally compact topological space, In that case we shall use the notation $\bf{R}$}. We will label the points of $\bf{R}$} as usual, so that we can write continuous functions, such as $x\to x^{2}$ or the involution $x\to-x$.
@@
    if and only if $\tau=id$ is the identity homeomorphism.
@@
    $z\overset{id}{\longrightarrow}z\qquad\qquad z\overset{\hat{\tau}}{\longrightarrow}\bar{z}\qquad\qquad z\overset{\sigma}{\longrightarrow}-z$
@@ 3. New $C^{*}$-algebras from Old Ones
Let $H$ be a real Hilbert space. Given a real $C^{*}$-algebra $A$, its complexification $A_{\mathbb{C}}=A+iA$ is a complex $C^{*}$-algebra:
  1. (1)
    $A_{\mathbb{C}}$ is a self-adjoint complex subalgebra.
@@
Suppose that $A$ and $B$ are real $C^{*}$-algebras. Then we may form their tensor product $A\otimes B$ by taking their algebraic tensor product over $\mathbb{R}$} and then completing that algebra in a suitable $C^{*}$-norm. The complication here is that in general, just as in the complex case, there is more than one “suitable” norm on the algebraic tensor product - there is a minimal norm, a maximal norm, and possibly many in between. The minimal norm has the very nice property of being concrete. The maximal norm has the very nice property of preserving exact sequences. We want both properties and hence we make a blanket assumption: whenever we write $A\otimes B$ then at least one of the two $C^{*}$-algebras has the property that its complexification is nuclear and exact. [^12] For example, any commutative complex $C^{*}$-algebra is nuclear and exact. Please see Blackadar [14] for definitions of nuclear and exact $C^{*}$-algebras as well as for other operator algebra terms that may seem mysterious. This does the job, because, using the minimal tensor product norm,
@@
    The complexification of a tensor product of two real $C^{*}$-algebras is isomorphic to the the tensor product of the complexification of the two real $C^{*}$-algebras, and
@@
    The same result holds in the category of real $C^{*}$-algebras.
@@
Assume first that $A_{n}$ is unital for all $n$ and that all of the connecting homomorphisms are unital. Assume also that $B$ is unital. The general case can be easily deduced from this. Let $A=\lim_{n\to\infty}A_{n}$ and let
@@
First of all, note that the sequence of maps
@@
which restrict to $f$ and $g$ on $A\otimes 1$ and $1\otimes B$ respectively.
@@
Here are some more examples:
@@ 4. Homotopy Constructions
    The bootstrap category [^14] First defined by Schochet in [73] (a certain subcategory of the previous example)
@@
    All commutative complex $C^{*}$-algebras and their homomorphisms151515which, by the Gelfand transform, corresponds to all compact spaces and continuous functions.
@@
    The real bootstrap category (a certain subcategory of the previous example as yet unexplored as far as we know)
@@
    All commutative real $C^{*}$-algebras and their homomorphisms161616which, as we have noted, does not correspond to all compact spaces and continuous functions but is larger!.
@@
The cylinder of $A$, denoted $IA$, is $\real([0,1],A)$ with canonical maps $p_{t}:IA\to A$ given by $p_{t}(\xi)=\xi(t)$.
@@
$SA=\{\xi\in IA\,|\,\xi(0)=\xi(1)=0\}\cong C_{o}^{{\mathbb{R}}}({\bf{R}},A)\cong C_{o}^{{\mathbb{R}}}({\bf{R}})\otimes A$
@@
$S^{n}A=C_{o}^{{\mathbb{R}}}({\bf{R}}^{n},A)\cong C_{o}^{{\mathbb{R}}}({\bf{R}}^{n})\otimes A,$
@@
$S^{n}A=C_{o}^{{\mathbb{R}}}({\bf{R}}^{n},C(X,{\mathbb{R}}))\cong C_{o}^{{\mathbb{R}}}({\bf{R}}^{n}\times X,{\mathbb{R}})$
@@
$(S^{n}A)^{+}\cong C(S^{n}\times X,{\mathbb{R}}).$
Taking $X=x_{0}$ a point, we have $C_{o}^{{\mathbb{R}}}(x_{0},{\mathbb{R}})={\mathbb{R}}$, and so
$S^{n}{\mathbb{R}}=C_{o}({\mathbb{R}}^{n},{\mathbb{R}})$
@@
$V^{n}A=V(V^{n-1}A)\,\cong\,C_{o}^{{\mathbb{R}}}(i{\bf{R}})^{\otimes n}\otimes A\,\cong\,C_{o}^{{\mathbb{R}}}({\bf{R}}^{n},\widetilde{\tau}^{\otimes n})\otimes A.$
@@
with $\hat{f}$ and $\hat{g}$ the evident projections.
The notion of “pullback” occurs in many different settings.191919and different names. Sometimes it is called the fibre product or the Cartesian square, for instance. Indeed, later in this work we will need to take the pullback of a vector bundle! It is a categorical definition and enjoys the following universal property: in the notation of the diagram above, suppose that one is given a real $C^{*}$-algebra $L$ and maps
@@
with $\pi\otimes 1$ a homotopy equivalence, so that there is a non-natural isomophism
$\mathcal{T}_{\mathbb{R}}\otimes A\cong\big(\mathcal{T}_{{\mathbb{R}}0}\otimes A\big)\oplus A.$
@@
The word ”cofibration” appears in classical algebraic topology. It has to do with a subspace $Y\subset X$ sitting nicely inside $X$. For example, the circle $S^{n}$ sits nicely inside the $n+1$ disk $D^{n+1}$ but the set of numbers $\{1,1/2,1/3,1/4,\dots\}$ does not sit nicely inside the closed unit interval.
We introduced the word ”cofibration” in [74] (1984) in the context of $C^{*}$-algebras so that the map
@@
is a cofibration. However, several authors have chosen to quote [74] but call this a ”fibration” (actually ”Schochet fibration”).  [^22] For example, Bunke [31] states the definition in much more elegant terms. He says that a map $A\to B$ is a Schochet fibration if the map $f_{*}:Hom(D,A)\longrightarrow Hom(D,B)$ is a Serre fibration of topological spaces for every $C^{*}$-algebra $D$. If $f:A\to B$ is a Schochet fibration with $J=Ker(f)$ then the associated short exact sequence $0\longrightarrow J\longrightarrow A\longrightarrow B\longrightarrow 0$ is a Schochet exact sequence. We will stick with ”cofibration.” The results in [74] are presented for complex $C^{*}$-algebras, but the parts having to do with cofibration apply to the real setting with no serious modification needed, so we will state facts but offer no proofs.
@@
    The map $p_{t}:IB\longrightarrow B$ is a cofibration for each $f\in[0,1]$. (Lemma 1.3)
@@
    then the pullback $f^{*}(p):C\to D$ is a cofibration. (Prop. 1.5)

### v3

# Real $K$-Theory for $C^{*}$-Algebras: just the facts January 13, 2026

@@ Abstract.
This monograph is intended to present the basic properties of $KO$-theory for real $C^{*}$-algebras and to explain its relationship with complex $K$-theory and with $KR$- theory. Whenever possible we will rely upon proofs in printed literature, particularly the work of Karoubi [52], Wood [89], Schröder [76], and more recent work of Boersema [15] and J. M. Rosenberg [70]. In addition, we shall explain how $KO$-theory is related to the Ten-Fold Way in physics and point out how some deeper features of $KO$-theory for operator algebras may provide powerful new tools there. Commutative real $C^{*}$-algebras not of the form $\real(X)$ will play a special role.
Unfortunately, there is no single reference for $KO$-theory for operator algebras that begins to compare with Blackadar’s wonderful exposition of complex $K$-theory [14]. This monograph is intended to provide a platform upon which mathematicians and mathematical physicists can rely in order to use these new tools in their research. As we are writing for a diverse audience of functional analysts, topologists, and physicists, we often present material well-known to one group of people and unfamiliar to another.
@@ 1. Introduction
First, a little history to put $KO$-theory into historical context. Let $R$ be an arbitrary ring. The group $K_{0}(R)$ was introduced by Alexander Grothendieck for arbitrary rings $R$ in 1957 as part of his revolution in algebraic geometry, as a way to formulate what is now known as the Grothendieck-Riemann-Roch theorem. It was based upon coherent sheaves on an algebraic variety. It was extended using modules over arbitrary rings and is difficult to compute in general. [^1] For everything you would want to know about the early history of algebraic $K$-theory, see Weibel [87]. The group $K_{1}(R)$ was defined in 1962, $K_{2}(R)$ in 1967, and $K_{n}(R)$ in 1970.
@@
As their paper was about to appear, Atiyah’s student R. Wood produced (1966)[89] a proof of the Periodicity Theorem along the lines suggested by ABS, where Clifford algebras play an essential role and where Bott Periodicity is a consequence of eight homotopy equivalences. He took a fundamental step forward, in that he stated the theorem in terms of the general linear groups of real and complex Banach algebras, and not just compact spaces, though his proof was rather sketchy. (See also Milnor [62]). This was followed by several detailed papers by Max Karoubi (cf. [45]). Karoubi generalized the Periodicity Theorem to Banach categories and Banach rings, proving periodicity in detail. Karoubi and Orlando Villamayor ([50], [51] ) explored its relationship to algebraic $K$-theory. A few years later Karoubi introduced Hermitian $K$-theory ([47], [49]), where he proved what he calls “A Fundamental Theorem” which is another road to Bott Periodicity (see the introduction to [47]). This theorem has important consequences nowadays in algebraic geometry and in quadratic forms over any field, not just the field of real numbers. Karoubi anticipated many of the results that we use in this monograph, In particular, the idea of desuspension was used to produce the negative algebraic $K$-theory groups developed by Bass.
@@
This monograph is designed to provide a basic source for those in mathematics and physics who wish to use $KO$-theory in their work. There is no such source at present: we mean to remedy that problem.
@@
Section 6 In Section 6 we introduce the homology theory $KO_{*}$ for real $C^{*}$-algebras, the main focus of this monograph. There are several distinct approaches to $KO$. We use what we believe to be the simplest approach, not relating much to algebraic $K$-theory, Hermitian $K$-theory, or other homotopy constructions. We show that $KO_{*}$ is a homology theory and introduce some of its relationships to $K_{*}$. Still no mention of Bott periodicity.
@@
Section 9 Section 9 introduces basic constructions in topological spaces such as the product, the join, the smash product, and the suspension. It establishes the fundamentally important identification $F(SX,Y)\cong F(X,\Omega Y)$. We will assume throughout the monograph that all topological spaces are compactly generated. We introduce $CW$-complexes briefly.
@@
Section 15 Section 15 is devoted to establishing the relationship between $KO^{*}$-theory for spaces and $KO_{*}$-theory for real $C^{*}$-algebras and similarly for $K^{*}$-theory for spaces and $K_{*}$-theory for complex $C^{*}$-algebras. We present a novel proof of Swan’s classical result relating the $KO_{*}$-theory of a real $C^{*}$-algebra $C^{\mathbb{R}}(X)$ and the topological $KO^{*}$-theory of the compact space $X$.
@@
Section 24 This section contains our treatment of $KO$-theory with coefficients. It is well-known that $H^{*}(X;G)$, the cohomology of a space with coefficients in a group $G$ is sometimes much more useful and certainly more tractable then when you take $G={\mathbb{Z}}$. We suspect that the same will be true for $KO$ and hence we present how to define it.
Section 25 This section presents a novel approach to the groups $K_{*}(A)$ and $KO_{*}(A)$, showing that each of the ten groups may be described concretely in terms of unitary elements of A and associated algebras such as $M_{k}(A)$.
What’s not in the monograph? We have not discussed the Künneth theorem in full generality, nor united $K$-theory, which combines $K$, KOKO and $KSc$ theories. We have not discussed some of the applications of $KO$-theory in differential geometry, as this requires quite a bit of background to appreciate. Nor have we discussed $KO$-homology theory for compact spaces. We have not discussed the Hopkins-Hovey isomorphism [43]
@@
This work arose in connection with the second author’s collaboration with Erik Akkermanns. We are most grateful to him and to his students, particularly to Nadav Orion, who have provided invaluable feedback and true insight in the development of this work. We are also very grateful to Jean Bellissard, Joachim Cuntz, Yigal Kamel, John Klein, Chris Phillips, Emil Prodan, Jonathan Rosenberg and Chuck Weibel, for providing us with various deep insights as this monograph developed. We especially want to acknowledge with great pleasure the extensive philosophical and detailed help that we have received from Max Karoubi, who, more than anyone, is responsible for the very existence of $KO$-theory for spaces and algebras as we know it.
@@ 2. Real $C^{*}$-algebras
Two points on notation. First, we have been confused over the years by various papers using $C(X)$, $K_{0}(A)$, and the like, without specifying whether they had real or complex versions in mind.777and we each plead guilty to this sin. Frequently authors deal with this by stating at the beginning that everything is complex or that everything is real. In this monograph both possibilities will arise. We will be more precise.
A second point. Usually the real numbers occur in this monograph as a real vector space or with their real $C^{*}$-algebra structure, and when that happens we shall denote them as $\mathbb{R}$}. Sometimes, though, we want to think of the real numbers simply as a locally compact topological space, In that case we shall use the notation $\bf{R}$}. We will label the points of $\bf{R}$} as usual, so that we can write continuous functions, such as $x\mapsto x^{2}$ or the involution $x\mapsto-x$.
@@
    if and only if $\tau=id$ is the identity homeomorphism. See Theorem 2.2 for a complete discussion of real commutative $C^{*}$-algebras.
@@
    $z\overset{id}{\longmapsto}z\qquad\qquad z\overset{\hat{\tau}}{\longmapsto}\bar{z}\qquad\qquad z\overset{\sigma}{\longmapsto}-z$
@@ 3. New $C^{*}$-algebras from Old Ones
Let $\mathcal{H}$ be a real Hilbert space. Given a real $C^{*}$-algebra $A$ of bounded operators on $\mathcal{H}$, its complexification is defined as the complex algebra
$A_{\mathbb{C}}=A+iA$
of bounded operators on $\mathcal{H}_{\mathbb{C}}$. Further,
@@
Note that any vector $\xi$ in $\mathcal{H}_{\mathbb{C}}$ can be written uniquely as $\xi=x+iy$ where $x,y\in\mathcal{H}$, and then
$\overline{\xi}=\overline{x+iy}=x-iy.$
@@
The definition and basic properties of the tensor product of two complex $C^{*}$-algebras are well-exposed in several sources, notably Blackadar [14]. For real $C^{*}$-algebras we rely upon Boersema [15].
@@
Suppose that $A$ and $B$ are real $C^{*}$-algebras. Then we may form their tensor product $A\otimes B$ by taking their algebraic tensor product over $\mathbb{R}$} and then completing that algebra in a suitable $C^{*}$-norm.  [^12] We will use the symbol $\otimes$ when it is clear from context that we are working with real or with complex $C^{*}$-algebras. If it is not clear then we shall use $\otimes_{\mathbb{R}}$ and $\otimes_{\mathbb{C}}$.
The complication here is that in general, just as in the complex case, there is more than one “suitable” norm on the algebraic tensor product - there is a minimal norm, a maximal norm, and possibly many in between. The minimal norm has the very nice property of being concrete. The maximal norm has the very nice property of preserving exact sequences. We want both properties and hence we make a blanket assumption: whenever we write $A\otimes B$ then at least one of the two $C^{*}$-algebras has the property that its complexification is nuclear and exact. [^13] For example, any commutative complex $C^{*}$-algebra is nuclear and exact. Please see Blackadar [14] for definitions of nuclear and exact $C^{*}$-algebras as well as for other operator algebra terms that may seem mysterious. This does the job, because, using the minimal tensor product norm,
@@
    The complexification of a tensor product of two real $C^{*}$-algebras is isomorphic to the tensor product of the complexification of the two real $C^{*}$-algebras, and
@@
    Similarly, if each $A_{n}$ is unital and the directed system preserves the identity then there is a natural map
[…]
    defined as a limit of the sequence of maps $B\hookrightarrow A_{n}\otimes B$ given by $b\mapsto 1\otimes b$.
  3. (3)
    The same results hold in the category of real $C^{*}$-algebras.
@@
[…]
@@
The sequence of maps
@@
which restricts to $f$ and $g$ on $A\otimes 1$ and $1\otimes B$ respectively.
@@
Here are some more examples, Let $A^{\otimes n}=A\otimes\dots\otimes A$ (n times). Then:
@@ 4. Homotopy Constructions
    The bootstrap category [^15] First defined by Schochet in [73]. This categroy is an important part of the hypotheses for the Künneth Formula and the Universal Coefficient Theorem (a certain subcategory of the previous example, which includes commutative $C^{*}$-algebras but is much larger),
@@
    All commutative complex $C^{*}$-algebras and their homomorphisms161616which, by the Gelfand transform, corresponds to all compact spaces and continuous functions.
@@
    The real bootstrap category (a certain subcategory of the previous example, analogous to the complex bootstrap category above, which includes all commutative real $C^{*}$-algebras)
@@
    The pre-bootstrap category, which consists of all real $C^{*}$-algebras $A$ whose complexifications $A_{\mathbb{C}}$ lie in the complex bootstrap category. This includes the real bootstrap category, by Corollary 4.3 of of [15].
  5. (5)
    All commutative real $C^{*}$-algebras and their homomorphisms171717which, as we have noted, does not correspond to all compact spaces and continuous functions but is larger!.
@@
The cylinder of $A$, denoted $IA$, is $C([0,1],A)$ with canonical maps $p_{t}:IA\to A$ given by $p_{t}(\xi)=\xi(t)$.
@@
$SA=\{\xi\in IA\,|\,\xi(0)=\xi(1)=0\}\cong C_{o}({\bf{R}},A)\cong C_{o}^{{\mathbb{R}}}({\bf{R}})\otimes A$
@@
$S^{n}A=C_{o}({\bf{R}}^{n},A)\cong C_{o}^{{\mathbb{R}}}({\bf{R}}^{n})\otimes A,$
@@
$S^{n}A=C_{o}^{{\mathbb{R}}}({\bf{R}}^{n},C(X,{\mathbb{R}}))\cong C_{o}^{{\mathbb{R}}}({\bf{R}}^{n}\times X)$
@@
$(S^{n}A)^{+}\cong\real(S^{n}\times X).$
Taking $X=x_{0}$ a point, we have $C_{o}^{{\mathbb{R}}}(x_{0})={\mathbb{R}}$, and so
$S^{n}{\mathbb{R}}=C_{o}^{{\mathbb{R}}}({\bf{R}}^{n})$
@@
$V^{n}A=V(V^{n-1}A)\,\cong\,C_{o}^{{\mathbb{R}}}(i{\bf{R}})^{\otimes n}\otimes A\,\cong\,C_{o}^{{\mathbb{R}}}({\bf{R}}^{n},\widetilde{\tau}^{\otimes n}\otimes A).$
where $\widetilde{\tau}(x)=-x$ for $x\in{\mathbb{R}}^{n}$ as in Example 8 in Section 2.
@@
and
$\hat{f}:D\to B\quad and\quad\hat{g}:D\to A$
are the evident projections.
The notion of “pullback” occurs in many different settings.202020and different names. Sometimes it is called the fibre product or the Cartesian square, for instance. Indeed, later in this monograph we will need to take the pullback of a vector bundle! It is a categorical definition and enjoys the following universal property: in the notation of the diagram above, suppose that one is given a real $C^{*}$-algebra $L$ and maps
@@
with $\pi\otimes 1$ a homotopy equivalence.
@@
The word “cofibration” appears in classical algebraic topology. It has to do with a subspace $Y\subset X$ sitting nicely inside $X$. For example, the circle $S^{n}$ sits nicely inside the $n+1$ disk $D^{n+1}$ but the set of numbers $\{1,1/2,1/3,1/4,\dots\}$ does not sit nicely inside the closed unit interval.
We introduced the word “cofibration” in [74] (1984) in the context of $C^{*}$-algebras so that the map
@@
is a cofibration. However, several authors have chosen to quote [74] but call this a “fibration” (actually “Schochet fibration”). [^23] For example, Bunke [31] states the definition in much more elegant terms. He says that a map $A\to B$ is a Schochet fibration if the map $f_{*}:Hom(D,A)\longrightarrow Hom(D,B)$ is a Serre fibration of topological spaces for every $C^{*}$-algebra $D$. If $f:A\to B$ is a Schochet fibration with $J=Ker(f)$ then the associated short exact sequence $0\longrightarrow J\longrightarrow A\longrightarrow B\longrightarrow 0$ is a Schochet exact sequence. We will stick with “cofibration.” The results in [74] are presented for complex $C^{*}$-algebras, but the parts having to do with cofibration apply to the real setting with no serious modification needed, so we will state facts but offer no proofs.
@@
    The map $p_{t}:IB\longrightarrow B$ is a cofibration for each $f\in[0,1]$. ([74] Lemma 1.3)
@@
    then the pullback $f^{*}(p):C\to D$ is a cofibration. ([74] Prop. 1.5)